OpenStudy (anonymous):

What is the sum? –25 + (–18)

OpenStudy (anonymous):

i give medal and fan!!!!!!!!!!!!!

OpenStudy (anonymous):

-45

OpenStudy (sidsiddhartha):

\[-25+(-18)=-25-18=-(25+18)\] do the rest :)

OpenStudy (anonymous):

@chosenmatt @cadence_sparkle<3 @cmcleod1 Direct answers are against the Code of Conduct. Why don't you follow @sidsiddhartha 's example, and lead the asker to the solution?

OpenStudy (chosenmatt):

sorry?

OpenStudy (anonymous):

Hello @hartnn . I assume you will take care of this problem.

OpenStudy (anonymous):

-43

OpenStudy (anonymous):

It's -43 because -25 + (-18) is the same as saying -25 -18. Then you add 25 + 18 and get 43, but don't forget to put the negative sign, so -43.

OpenStudy (anonymous):

Wow... So much direct answers.

OpenStudy (sidsiddhartha):

here some basic algebra \[a+b=a+b\\a+(-b)=a-b\\a-(-b)=a+b\\-a-b=-(a+b)\\-a-(-b)=-a+b\]

OpenStudy (sidsiddhartha):

hope this helps :)

OpenStudy (anonymous):

thanks it helped i am ready to pass